# On the integrability of extended test body dynamics around black holes

### v2

@@ Abstract
In general relativity, the motion of a test extended body is influenced by its proper rotation, or spin. We present a covariant and physically self-consistent Hamiltonian framework to study this motion, up to quadratic order in the body’s spin, including a spin-induced quadrupole, and in an arbitrary background spacetime. The choice of spin supplementary condition and degeneracies associated with local Lorentz invariance are treated rigorously with adapted tools from Hamiltonian mechanics. Applying the formalism to a background space-time described by the Kerr metric, we prove that the motion of any test compact object around a rotating black hole defines an integrable Hamiltonian system to linear order in the body’s spin. Moreover, this integrability still holds at quadratic order in spin when the compact object has the deformability expected for an isolated black hole. Our analytical results shed light on longstanding numerical conjectures regarding spin-induced chaos in the motion of asymmetric compact binaries, and provides a powerful framework to improve current gravitational waveform modelling to account for spin-induced extended body effects.
Introduction.— The motion of a celestial body in a given gravitational field is, arguably, the oldest and most fundamental problem in theoretical astrophysics. Although it generally resists simple, analytical solutions – mostly due to the complex internal composition of celestial bodies – the bulk motion of such a body adheres to conservation laws for linear and angular momentum. The microscopic details enter the dynamics at finer levels of description. This remarkable feature is consequential to the universality of gravitation, the core of Einstein’s theory of general relativity. The natural setup to understand these properties is through multipolar expansions of a body Mathisson (1937); Dixon (1964); Harte (2015); Mathews et al. (2022). In general relativity, these schemes assert that, aside from dissipation, the motion of an extended body in a given spacetime mirrors the worldline of a point object endowed with multipoles. At monopolar order, the body is replaced by a point mass whose worldline is a geodesic of the background spacetime: A mere rephrasing of Einstein’s geodesic principle Einstein (1914). However, the latter is only an approximation: At dipolar order, the body’s proper rotation (hereafter spin) couples to spacetime’s curvature, thus deviating its worldline from a geodesic. Remarkably, this dipolar, non-geodesic motion is still universal: The body’s momenta obey purely kinematical evolution equations along a worldline that is completely determined by the background geometry Harte (2015). Universality disappears at quadrupolar order, with body-specific multipoles entering the description and dynamically driving the evolution equations via self-forces and torques Harte and Dwyer (2023).
@@
Naturally, the integrability of Kerr geodesics has prompted many to question the integrability around black holes at higher multipolar order, where, in particular, the body’s spin breaks geodesic motion and Carter’s constant is not conserved any more. To speak of integrability, one requires a Hamiltonian formulation of the non-dynamics beyond geodesics. Several have been proposed Barausse et al. (2009); Steinhoff (2011); d’Ambrosi et al. (2015); Vines et al. (2016); Witzany et al. (2019) for other objectives than showing integrability. However, these frameworks either lack covariance, have too many dimensions or implement the (crucial) spin supplementary condition heuristically. These features are not intrinsically problematic, but limit one’s ability to apply some tools of classical Hamiltonian theory, thereby resulting in debatable conclusions about integrability (or lack thereof) Kunst et al. (2016); Witzany (2019); Compère and Druart (2022); Grant and Moxon (2023). Supported by numerical studies revealing signatures of chaotic motion Kunst et al. (2016); Lukes-Gerakopoulos (2018); Witzany et al. (2019), this has led to the current conjecture that a body’s spin should generally break integrability around rotating black holes, even at linear order in spin.
@@
In this work, we propose to go beyond the pre-existing numerical and/or conjectural results, and answer analytically whether the body’s spin breaks Carter’s geodesic integrability in a Kerr background. This is done in three steps. First, we describe the quadratic-in-spin dynamics of an extended test body in any spacetime. Our framework is at once covariant, non-degenerate, 5-dimensional and does not require any heuristic or external ad-hoc parameter-fixing. Second, when the framework is applied to a background describing a Kerr black hole, the resulting system is proven to be integrable, owing to the existence of enough independent integrals of motion for any compact object. Third, still in a Kerr background, we show that quadratic-in-spin terms (from both dipole effects and a spin-induced quadrupole) do not break integrability when the test object has the quadrupole structure expected for an isolated black hole.
@@
Geometrical units $G=c=1$ are used throughout, and our conventions for Lorentzian and symplectic geometry follow Wald (1984) and Arnold et al. (2006), respectively. Extended discussions and detailed calculations are presented in Ramond (2024) and Ramond and Druart (2024) for linear and quadratic order in spin, respectively.
@@
Quadrupolar approximation.— We consider a fixed background spacetime $(\mathscr{E},g_{ab})$ where $\mathscr{E}$ is a 4-dimensional (4D) manifold covered with coordinates $x^{\alpha}$, and $g_{ab}$ is a metric tensor on $\mathscr{E}$. Following classical multipolar expansion schemes Dixon (1974); Harte (2015), the motion of an extended body in $\mathscr{E}$ can be effectively described by the trajectory of a representative point particle endowed with a number of multipoles. Let $\mathscr{L}\subset\mathscr{E}$ be that particle’s worldline, $\tau$ its proper time and $u^{a}$ its four-velocity, normalized as $g_{ab}u^{a}u^{b}=-1$. The multipolar expansion accounts for both translational and rotational degrees of freedom of the body, encoded into a momentum 1-form $p_{a}$ and an antisymmetric spin tensor $S^{ab}$. Evolution equations for $(p_{a},S^{ab})$ are known as the MPTD equations Mathisson (1937); Papapetrou (1951); Tulczyjew (1959); Dixon (1974); Harte (2012)
@@
$J^{abcd}:=\frac{3\kappa}{\mu}\bar{p}^{[a}\Theta^{a][b}\bar{p}^{d]}\,,\,\,\text{with}\,\,\begin{cases}\,\Theta^{ab}:=g_{cd}S^{ac}S^{db}\,,\\ \,\,\bar{p}^{a}:=p_{a}/\mu\,,\end{cases}$ (2)
and where $\kappa$ is a dimensionless coupling parameter that encodes the deformability of the extended object. Its value is expected to be $1$ for an isolated Kerr black hole Binnington and Poisson (2009), and $\kappa>1$ for other compact object, larger values of $\kappa$ indicating “stiffer” equations of state Laarakkers and Poisson (1999); Boshkayev et al. (2012); Uchikata et al. (2016). The system (1)–(2) suffers from two drawbacks. First, it is physically inconsistent: Cubic-in-$S^{ab}$ terms are known to arise at the octupolar order Marsat (2015) and contribute to the right-hand sides of Eqs. (1). Therefore, any spin-induced quadrupolar model is only self-consistent at quadratic order in $S^{ab}$. Second, the system is mathematically ill-posed: For a given spacetime geometry, it is equivalent to 10 ordinary differential equations for 13 unknowns in $(u^{a},p_{a},S^{ab})$. Both issues are solved by choosing a spin supplementary condition (SSC) Costa and Natário (2015). For well-motivated reasons exposed in Ramond (2024), we choose the Tulczyjew SSC:
@@
Physically, Eq. (3) is interpreted as the vanishing of the body’s mass dipole $\bar{C}^{b}:=\bar{p}_{a}S^{ab}$ as measured by an observer of four-velocity $\bar{p}^{a}$. Combining Eqs. (1) with the covariant derivative of (3) leads to the momentum-velocity relation Gralla et al. (2010)
$u^{a}=\bar{p}^{a}+\frac{1}{2\mu}R_{abcd}S^{ab}\bar{p}^{d}S^{ac}-\frac{2}{3\mu}J^{[a}_{\phantom{a}cde}R^{b]cde}\bar{p}_{b}\,,$ (4)
where the result has been expanded consistently to quadratic order in $S^{ab}$, for self-consistency of the quadrupolar approximation. This will be always done subsequently. Combining Eqs. (1)–(4), it follows that that $S_{\star}$ vanishes identically Witzany et al. (2019), and that both $S_{\circ}$ and the effective mass $\tilde{\mu}:=\mu-\frac{1}{6}R_{abcd}J^{abcd}$ are conserved along $\mathscr{L}$ at quadratic order in spin Dixon (1970); Marsat (2015); Compère et al. (2023). Projecting Eqs. (1)–(4) onto the coordinate basis $(\partial_{\alpha})^{a}$ and characterizing the worldline in the form $u^{\alpha}=\mathrm{d}x^{\alpha}/\mathrm{d}\tau$ leads to 14 first order, ordinary differential equations for 14 unknowns $x^{\alpha}$, $p_{\alpha}$ and $S^{\alpha\beta}$.
Formulation as a Poisson system.— We now wish to turn system (1)–(4) into a Hamiltonian system [^1] Our framework is that of symplectic geometry and not that of Hamiltonian built from the (possibly singular) Legendre transform of some Lagrangian, as is customary. in the general sense, i.e., as defined from the three following ingredients: (i) a phase space $\mathcal{M}$, defined as a $N$-dimensional manifold endowed with $N$ coordinates $y:=(y^{1},\ldots,y^{N})\in\mathbb{R}^{N}$ ; (ii) a Poisson structure $\Lambda$, defined as a skew-symmetric $N\times N$ matrix whose entries, denoted $\Lambda^{ij}(y)$, satisfy the Jacobi identity $\Lambda^{\ell(i}\partial_{\ell}\Lambda^{jk)}=0$, with $\partial_{\ell}:=\partial/\partial y^{\ell}$ ; (iii) a Hamiltonian $H$, defined as a scalar field $\mathcal{M}\rightarrow\mathbb{R}$. The triplet $(\mathcal{M},\Lambda,H)$ is called a Poisson system. If, for all $y\in\mathcal{M}$, the matrix $\Lambda(y)$ has maximal rank, one speaks of a symplectic structure. If not, it is said to be degenerate. Regardless of its degeneracy, the Poisson structure $\Lambda$ defines a Poisson bracket via:
@@
where $\delta^{\alpha}_{\beta}$ is the 4D Kronecker symbol. To our knowledge, there does not exist a Hamiltonian in the literature that, when varied with respect to the brackets (6), generates the covariant, quadrupolar MPTD equations (1)–(4), (see, however, Vines et al. (2016)). We found such a Hamiltonian by exploiting the conservation of the effective mass $\tilde{\mu}$, which is the sum of the dynamical mass $\mu$ (conserved at linear order in spin, cf. Ramond (2024)) and a quadrupolar correction. Based on these observations, we construct the natural Hamiltonian $H:=H_{g}+H_{d}+H_{q}$ as that of geodesic motion ($H_{g}$) with quadratic-in-spin terms due to the dipolar $(H_{d})$ and quadrupolar $(H_{q})$ sector. Explicitly:
$H:=\frac{1}{2}g^{\alpha\beta}p_{\alpha}p_{\beta}+\bar{p}^{\alpha}R_{\alpha\beta\gamma\delta}S^{\beta\gamma}\bar{C}^{\delta}-\frac{\kappa}{2}R_{\alpha\beta\gamma\delta}\Theta^{\beta\gamma}\bar{p}^{\alpha}\bar{p}^{\delta}\,.$ (7)
Notice that $H$ has no linear-in-$S^{ab}$ piece: such terms will come from the geometry (6) alone, which is expected from the universality of linear-in-spin motion Harte (2012). In Eqs. (6) and (7), it is important to appreciate that the phase space coordinates $(x^{\alpha},p_{\alpha},S^{\alpha\beta})$ can be non-trivially hidden in each term. Taking all these dependences into account and replacing $F$ by each coordinate in Hamilton’s equation gives back Eqs. (1) and (4), provided that $\lambda=\tau/\mu=:\bar{\tau}$. As the Hamiltonian (7) is autonomous (independent of $\bar{\tau}$), it is conserved along any solution to Hamilton’s equation. Indeed, under Eq. (3), $H=-\tilde{\mu}^{2}/2$, which, as expected, must be conserved along $\mathscr{L}$.
Degeneracy of the system.— The 14D Poisson manifold $(\mathcal{M},\{,\})$ is degenerate Vines et al. (2016). These degeneracies become obvious when we switch to the variables $(x^{\alpha},\pi_{\alpha},S^{I},D^{I})$, where
$\pi_{\alpha}:=p_{\alpha}+\tfrac{1}{2}\omega_{\alpha BC}S^{BC}\,,\,\,S^{I}:=\tfrac{1}{2}\varepsilon^{I}_{\phantom{I}JK}S^{JK}\,,\,\,D^{I}:=S^{0I}\,,$ (8)
@@
Symplectic formulation.— The classical notion of Liouville-Arnold integrability requires a symplectic (non-degenerate) formulation of the dynamics. Weinstein splitting theorem’s for Poisson Manifolds Vaisman (2012) shows that the 14D Poisson manifold $(\mathcal{M},\Lambda)$ is foliated by 12D, symplectic sub-manifolds known as symplectic leaves, corresponding to the level sets of the Casimirs $(\mathcal{C}_{1},\mathcal{C}_{2})$. Let $\mathcal{N}$ be such a leaf, where the Casimirs have physical value $(S_{\circ}^{2},S_{\star}^{2})$. On $\mathcal{N}$, there exists a natural non-degenerate Poisson structure inherited from that of $\mathcal{M}$. Symplecticity means that it is possible, at least locally, to endow $\mathcal{N}$ with 12 canonical coordinates, i.e., 6 pairs $(q^{i},\pi_{q_{i}})_{i\in\{1,\ldots,6\}}$ for which the brackets are $\{q^{i},\pi_{q_{j}}\}=\delta^{i}_{j}$. These coordinates are given explicitly in Ramond (2024) but are not required for our present purposes. The Hamiltonian on $\mathcal{N}$ is obtained by inserting the new coordinates (8) in the Hamiltonian (7) on $\mathcal{M}$. One finds
@@
where $Q_{ABCD}:=\tfrac{1}{8}\omega^{\alpha}_{\phantom{\alpha}AB}\omega_{\alpha CD}+R_{EABD}\bar{\pi}^{E}\bar{\pi}_{C}-\tfrac{\kappa}{2}R_{EADF}\bar{\pi}^{E}\bar{\pi}^{F}\eta_{BC}$ depends only on $(x^{\alpha},\pi_{\alpha})$ and should be set to zero to recover the linear-in-spin dynamics studied in Ramond (2024); Ramond and Isoyama (2024). We note that the quantities $\bar{\pi}^{A}:=\eta^{AB}(e_{B})^{\alpha}\pi_{\alpha}/\mu$ depend on the coordinates $x^{\alpha}$ (via $(e_{B})^{\alpha}$ and $\mu$) and $\pi_{\alpha}$ (explicitly and in $\mu$), while $S^{AB}$ are simple functions of $(S^{I},D^{I})$ hidden through $S^{0I}=D^{I}$ and $S^{IJ}=\varepsilon^{IJ}_{\phantom{IJ}K}S^{K}$, cf. Eq. (8). The Hamiltonian (12) on the 12D leaves $\mathcal{N}$, along with the 12 canonical coordinates given in Ramond (2024) and canonical Poisson brackets, define a 12D symplectic Hamiltonian system. This system, however, is not physically satisfactory: it contains unphysical trajectories and hosts spurious chaotic regions unrelated to physical effects. This is a crucial point of our analysis, which we now detail.
@@
Hamiltonian status of the SSC.— The phase space $\mathcal{N}$ is 12-dimensional, even though the physics at hand, with the three independent SSC constraints (3) and the normalization condition $u_{\alpha}u^{\alpha}=-1$, present only $14-3-1=10$ independent unknowns. In addition, $\mathcal{N}$ is filled with trajectories that are unphysical, namely those that do not satisfy the SSC (3). These issues come from a subtlety of the MPTD system scarcely noted and worth discussing. Equations (1)-(4) are a consequence of the SSC (3), but does not imply it back: Eqs. (1)-(4) only imply $\nabla_{u}C^{a}=0$. Parallel transport along $\mathscr{L}$ being linear, if $C^{a}=0$ at some point on $\mathscr{L}$, then $\nabla_{u}C^{a}=0$ implies $C^{a}=0$ at each point of $\mathscr{L}$. In other words, when working at the level of the evolution equations along $\mathscr{L}$ with some initial conditions, this issue is harmless. However, in a symplectic framework, there is no way to “choose” such initial condition. The phase space $\mathcal{M}$ and its leaves $\mathcal{N}$ are swarmed with trajectories that satisfy Hamilton’s equations (4) but not the SSC (3), making them physically irrelevant.

### v3

@@ Abstract
In general relativity, the motion of an extended test body is influenced by its proper rotation, or spin. We present a covariant and physically self-consistent Hamiltonian framework to study this motion, up to quadratic order in the body’s spin, including a spin-induced quadrupole, and in an arbitrary background spacetime. The choice of spin supplementary condition and degeneracies associated with local Lorentz invariance are treated rigorously with adapted tools from Hamiltonian mechanics. Applying the formalism to a background space-time described by the Kerr metric, we prove that the motion of any test compact object around a rotating black hole defines an integrable Hamiltonian system to linear order in the body’s spin. Moreover, this integrability still holds at quadratic order in spin when the compact object has the deformability expected for an isolated black hole. By exploiting the unique symmetries at play in black hole binaries, our analytical results clarify longstanding numerical conjectures regarding spin-induced chaos in the motion of asymmetric compact binaries, and may provide a powerful framework to improve current gravitational waveform modelling.
Introduction.— The motion of a celestial body in a given gravitational field is, arguably, the oldest and most fundamental problem in theoretical astrophysics. Although it generally resists simple, analytical solutions – mostly due to the complex internal composition of celestial bodies – the bulk motion of such a body adheres to conservation laws for linear and angular momentum. The microscopic details enter the dynamics at finer levels of description. This remarkable feature is consequential to the universality of gravitation, the core of Einstein’s theory of general relativity. The natural setup to understand these properties is through general relativistic multipolar expansions of a body Mathisson (1937); Dixon (1964); Harte (2015); Mathews et al. (2022). These schemes assert that, aside from dissipation, the motion of an extended body in a given spacetime mirrors the worldline of a point object endowed with multipoles. At monopolar order, the body is replaced by a point mass whose worldline is a geodesic of the background spacetime: A mere rephrasing of Einstein’s geodesic principle Einstein (1914). At dipolar order, the body’s proper rotation (hereafter spin) couples to spacetime’s curvature, thus deviating its worldline from a geodesic. Remarkably, this dipolar, non-geodesic motion is still universal: The body’s momenta obey purely kinematical evolution equations along a worldline that is completely determined by the background geometry Harte (2015). Universality disappears at quadrupolar order, with body-specific multipoles entering the description and dynamically driving the evolution equations via self-forces and torques Harte and Dwyer (2023).
@@
Naturally, the integrability of Kerr geodesics has prompted many to question the integrability around black holes at higher multipolar order, where, in particular, the body’s spin breaks geodesic motion and Carter’s constant is not conserved anymore. To speak of integrability, one requires a Hamiltonian formulation of the dynamics beyond geodesics. Several have been proposed Barausse et al. (2009); Steinhoff (2011); d’Ambrosi et al. (2015); Vines et al. (2016); Witzany et al. (2019) for other objectives than showing integrability. However, most of these frameworks are limited to linear-in-spin dynamics, and either lack general covariance, work with degenerate phase spaces with too many dimensions and/or implement the (crucial) spin supplementary condition heuristically. While not intrinsically problematic for specific goals, these aforementioned features limit one’s ability to apply standard tools of classical Hamiltonian theory, thereby resulting in debatable conclusions about integrability (or lack thereof) Kunst et al. (2016); Witzany (2019); Compère and Druart (2022); Grant and Moxon (2023). Supported by numerical studies of the revealing signatures of chaotic motion Suzuki and Maeda (1997, 1999); Hartl (2003); Hartl2 (2003); Kao and Cho (2005); Han (2008); Kunst et al. (2016); Lukes-Gerakopoulos (2018); Witzany et al. (2019), this has led to the current conjecture that a body’s spin should generally break integrability around rotating black holes, even at linear order in spin.
In this work, we propose to go beyond the pre-existing numerical and/or conjectural results, and answer analytically whether the body’s spin breaks Carter’s geodesic integrability in a Kerr background. Reaching this goal up to quadratic order in the body’s spin, we found, required to step aside from common practices, such as the use of (i) Marck tetrads Marck (1983); Van de Meent (2020) (limited to the linear-in-spin dynamics and requiring the introduction of extra fiducial parameters in the system Witzany (2019)); (ii) 3+1 spacetime decomposition and the Newton-Wigner supplementary condition Kunst et al. (2016); Vines et al. (2016) (which break general covariance and lead to a pure-constraint Hamiltonian via singular Legendre transform); and (iii) the use of too large a phase space and/or with degenerate Poisson structure on it Compère and Druart (2022) (which leads to a non-integrable, but also non-physical system, at least in Kerr). All these aforementioned methods, and their (sometimes) unpractical consequences, are absent from our framework. On the contrary, everything that follows relies solely on covariant calculations and tools from Poisson geometry Vaisman (2012).
@@
Our results can be summarized as the outcomes of a three-steps process. First, we describe the quadratic-in-spin dynamics of an extended test body in any spacetime as a Hamiltonian system that is at once covariant, non-degenerate, on a 10-dimensional phase space and does not require any heuristic or external ad-hoc parameter-fixing. Second, when the framework is applied to a background describing a Kerr black hole, the resulting system is proven to be integrable at linear order in spin, owing to the existence of enough independent integrals of motion for any test body. Third, still in a Kerr background, we show that quadratic-in-spin terms (from both dipole effects and a spin-induced quadrupole) do not break integrability when the test object has the quadrupole structure expected for an isolated black hole. Possible applications of the results are then discussed.
@@
Geometrical units $G=c=1$ are used throughout, and our conventions for Lorentzian and Poisson geometry follow Wald (1984) and Arnold et al. (2006), respectively.
@@
Quadrupolar dynamics.— We consider a fixed background spacetime $(\mathscr{E},g_{ab})$ where $\mathscr{E}$ is a 4-dimensional (4D) manifold covered with coordinates $x^{\alpha}$, and $g_{ab}$ is a metric tensor on $\mathscr{E}$. Following classical multipolar expansion schemes Dixon (1974); Harte (2015), the motion of an extended body in $\mathscr{E}$ can be effectively described by the trajectory of a representative point particle endowed with a number of multipoles. Let $\mathscr{L}\subset\mathscr{E}$ be that particle’s worldline, $\tau$ its proper time and $u^{a}$ its four-velocity, normalized as $g_{ab}u^{a}u^{b}=-1$. The multipolar expansion accounts for both translational and rotational degrees of freedom of the body, encoded into a momentum 1-form $p_{a}$ and an antisymmetric spin tensor $S^{ab}$. The evolution equations for $(p_{a},S^{ab})$ along $\mathscr{L}$ are known as the Mathisson-Papapetrou-Tulczyjew-Dixon (MPTD) equations Mathisson (1937); Papapetrou (1951); Tulczyjew (1959); Dixon (1974); Harte (2012)
@@
$J^{abcd}:=\frac{3\kappa}{\mu}\bar{p}^{[a}\Theta^{b][c}\bar{p}^{d]}\,,\,\,\text{with}\,\,\begin{cases}\,\Theta^{ab}:=g_{cd}S^{ac}S^{db}\,,\\ \,\,\bar{p}^{a}:=p^{a}/\mu\,,\end{cases}$ (2)
and where $\kappa$ is a dimensionless coupling parameter that encodes the deformability of the extended object. Its value is expected to be $1$ for an isolated Kerr black hole Binnington and Poisson (2009), and $\kappa>1$ for other compact object, larger values of $\kappa$ indicating “stiffer” equations of state Laarakkers and Poisson (1999); Boshkayev et al. (2012); Uchikata et al. (2016). The system (1)–(2) suffers from two drawbacks. First, it is physically inconsistent: Cubic-in-$S^{ab}$ terms are known to arise at the octupolar order Marsat (2015) and contribute to the right-hand sides of Eqs. (1). Therefore, any spin-induced quadrupolar model is only self-consistent at quadratic order in $S^{ab}$. Second, the system is mathematically ill-posed: For a given spacetime geometry, it is equivalent to 10 ordinary differential equations for 13 unknowns in $(u^{a},p_{a},S^{ab})$. These issues are solved in two steps, as follows. First, as is customary, we start by closing the differential system (1) with four algebraic equations known as a spin supplementary condition (SSC) Costa and Natário (2015). For well-motivated reasons exposed in Ramond (2024), we choose the Tulczyjew SSC:
@@
Physically, Eq. (3) is interpreted as the vanishing of the body’s mass dipole $\bar{C}^{b}:=\bar{p}_{a}S^{ab}$ as measured by an observer of four-velocity $\bar{p}^{a}$. Second, we combine Eqs. (1) with the covariant derivative of (3) to obtain a momentum-velocity relation Gralla et al. (2010)
$u^{a}=\bar{p}^{a}+\frac{1}{2\mu^{2}}R_{ebcd}S^{eb}\bar{p}^{d}S^{ac}+\frac{4}{3\mu}J^{[a}_{\phantom{a}cde}R^{b]cde}\bar{p}_{b}\,,$ (4)
where the result has been expanded consistently to quadratic order in $S^{ab}$, for self-consistency of the quadrupolar approximation. This will be always done subsequently. Combining Eqs. (1)–(4), it follows that $S_{\star}$ vanishes identically Witzany et al. (2019), and that both $S_{\circ}$ and the effective mass $\tilde{\mu}:=\mu-\frac{1}{6}R_{abcd}J^{abcd}$ are conserved along $\mathscr{L}$ at quadratic order in spin Dixon (1970); Marsat (2015); Compère et al. (2023). Projecting Eqs. (1)–(4) onto the coordinate basis $(\partial_{\alpha})^{a}$ and characterizing the worldline in the form $u^{\alpha}=\mathrm{d}x^{\alpha}/\mathrm{d}\tau$ leads to 14 first order, ordinary differential equations for 14 unknowns $x^{\alpha}$, $p_{\alpha}$ and $S^{\alpha\beta}$.
Formulation as a Poisson system.— We now wish to turn system (1)–(4) into a Hamiltonian system in the general sense, i.e., as defined from the three following ingredients: (i) a phase space $\mathcal{M}$, defined as a $N$-dimensional manifold endowed with $N$ coordinates $y:=(y^{1},\ldots,y^{N})\in\mathbb{R}^{N}$ ; (ii) a Poisson structure $\Lambda$, defined as a skew-symmetric $N\times N$ matrix whose entries, denoted $\Lambda^{ij}(y)$, satisfy the Jacobi identity $\Lambda^{\ell(i}\partial_{\ell}\Lambda^{jk)}=0$, with $\partial_{\ell}:=\partial/\partial y^{\ell}$ ; (iii) a Hamiltonian $H$, defined as a scalar field $\mathcal{M}\rightarrow\mathbb{R}$. The triplet $(\mathcal{M},\Lambda,H)$ is called a Poisson system. If, for all $y\in\mathcal{M}$, the matrix $\Lambda(y)$ has maximal rank, one speaks of a symplectic structure. If not, it is said to be degenerate. Regardless of its degeneracy, the Poisson structure $\Lambda$ defines a Poisson bracket via:
@@
where $\delta^{\alpha}_{\beta}$ is the 4D Kronecker symbol. To our knowledge, there does not exist a Hamiltonian in the literature that, when varied with respect to the brackets (6), generates the covariant, background-independent, quadrupolar MPTD equations (1)–(4). We found such a Hamiltonian by exploiting the conservation of the effective mass $\tilde{\mu}$, which is the sum of the dynamical mass $\mu$ (conserved at linear order in spin, cf. Ramond (2024)) and a quadrupolar correction. Based on these observations, we construct the natural Hamiltonian $H:=H_{g}+H_{d}+H_{q}$ as that of geodesic motion ($H_{g}$) with quadratic-in-spin terms due to the dipolar $(H_{d})$ and quadrupolar $(H_{q})$ sector. Explicitly:
$H:=\frac{1}{2}g^{\alpha\beta}p_{\alpha}p_{\beta}+R_{\alpha\beta\gamma\delta}\bar{p}^{\alpha}S^{\beta\gamma}\bar{C}^{\delta}+\frac{\kappa}{2}R_{\alpha\beta\gamma\delta}\bar{p}^{\alpha}\Theta^{\beta\gamma}\bar{p}^{\delta}\,.$ (7)
Notice that $H$ has no linear-in-$S^{ab}$ piece: such terms will come from the geometry (6) alone, which is expected from the universality of linear-in-spin motion Harte (2012). In Eqs. (6) and (7), it is important to appreciate that the phase space coordinates $(x^{\alpha},p_{\alpha},S^{\alpha\beta})$ can be non-trivially hidden in each term. Taking all these dependencies into account and replacing $F$ by each coordinate in Hamilton’s equation gives back Eqs. (1) and (4), provided that $\lambda=\tau/\mu=:\bar{\tau}$. This is a non-trivial calculation that ensures the validity of the whole Hamiltonian framework. As the Hamiltonian (7) is autonomous (independent of $\bar{\tau}$), it is conserved along any solution to Hamilton’s equation. Indeed, under Eq. (3), $H=-\tilde{\mu}^{2}/2$, which, as expected, must be conserved along $\mathscr{L}$.
Degeneracy of the system.— The 14D Poisson manifold $(\mathcal{M},\{,\})$ is degenerate Witzany et al. (2019). These degeneracies become obvious when we switch to the variables $(x^{\alpha},\pi_{\alpha},S^{I},D^{I})$, where
$\pi_{\alpha}:=p_{\alpha}-\tfrac{1}{2}\omega_{\alpha BC}S^{BC}\,,\,\,S^{I}:=\tfrac{1}{2}\varepsilon^{I}_{\phantom{I}JK}S^{JK}\,,\,\,D^{I}:=S^{0I}\,,$ (8)
@@
Symplectic formulation.— The classical notion of Liouville-Arnold integrability requires a symplectic (non-degenerate) formulation of the dynamics. Weinstein splitting theorem’s for Poisson manifolds Vaisman (2012) shows that the 14D Poisson manifold $(\mathcal{M},\Lambda)$ is foliated by 12D, symplectic sub-manifolds known as symplectic leaves, corresponding to the level sets of the Casimirs $(\mathcal{C}_{1},\mathcal{C}_{2})$. Let $\mathcal{N}$ be such a leaf, where the Casimirs have physical value $(S_{\circ}^{2},S_{\star}^{2})$. On $\mathcal{N}$, there exists a natural non-degenerate Poisson structure inherited from that of $\mathcal{M}$. Symplecticity means that it is possible, at least locally, to endow $\mathcal{N}$ with 12 canonical coordinates, i.e., 6 pairs $(q^{i},\pi_{q_{i}})_{i\in\{1,\ldots,6\}}$ for which the brackets are $\{q^{i},\pi_{q_{j}}\}=\delta^{i}_{j}$. These coordinates can be built explicitly Ramond (2024) but are not required for our present purposes. The Hamiltonian on $\mathcal{N}$ is obtained by inserting the new coordinates (8) in the Hamiltonian (7) on $\mathcal{M}$. One finds
@@
where $Q_{ABCD}:=\tfrac{1}{8}\omega^{\alpha}_{\phantom{\alpha}AB}\omega_{\alpha CD}+R_{EABD}\bar{\pi}^{E}\bar{\pi}_{C}+\tfrac{\kappa}{2}R_{EADF}\bar{\pi}^{E}\bar{\pi}^{F}\eta_{BC}$ depends only on $(x^{\alpha},\pi_{\alpha})$ and should be set to zero to recover the linear-in-spin dynamics studied in Ramond (2024); Ramond and Isoyama (2024). We note that the quantities $\bar{\pi}^{A}:=\eta^{AB}(e_{B})^{\alpha}\pi_{\alpha}/\mu$ depend on the coordinates $x^{\alpha}$ (via $(e_{B})^{\alpha}$ and $\mu$) and $\pi_{\alpha}$ (explicitly and in $\mu$), while $S^{AB}$ are simple functions of $(S^{I},D^{I})$ hidden through $S^{0I}=D^{I}$ and $S^{IJ}=\varepsilon^{IJ}_{\phantom{IJ}K}S^{K}$, cf. Eq. (8). The Hamiltonian (12) on the 12D leaves $\mathcal{N}$, along with the 12 canonical coordinates given in Ramond (2024) and canonical Poisson brackets, define a 12D symplectic Hamiltonian system. This system, however, is not physically satisfactory: it contains unphysical trajectories and hosts spurious chaotic regions unrelated to physical effects. This is a crucial point of our analysis, which we now detail.
@@
Hamiltonian status of the SSC.— The phase space $\mathcal{N}$ is 12-dimensional, even though the physics at hand, with the three independent SSC constraints (3) and the normalization condition $u_{\alpha}u^{\alpha}=-1$, present only $14-3-1=10$ independent unknowns. In addition, $\mathcal{N}$ is filled with trajectories that are non-physical, namely those that do not satisfy the SSC (3). These issues come from a subtlety of the MPTD system scarcely noted and worth discussing. Equations (1)-(4) are a consequence of the SSC (3), but does not imply it back: Eqs. (1)-(4) only imply $\nabla_{u}C^{a}=0$. Parallel transport along $\mathscr{L}$ being linear, if $C^{a}=0$ at some point on $\mathscr{L}$, then $\nabla_{u}C^{a}=0$ implies $C^{a}=0$ at each point of $\mathscr{L}$. In other words, when working at the level of the evolution equations along $\mathscr{L}$ with some initial conditions, this issue is harmless. However, in a symplectic framework, there is no way to “choose” such initial condition. The phase space $\mathcal{M}$ and its leaves $\mathcal{N}$ are swarmed with trajectories that satisfy Hamilton’s equations (4) but not the SSC (3), making them physically irrelevant.
What one must do is restrict the analysis to the sub-manifold $\mathcal{T}\subset\mathcal{M}$ where the $p_{\alpha}S^{\alpha\beta}=0$ holds identically. This sub-manifold has two important properties. First, it has co-dimension 3 in $\mathcal{M}$, as (3) only has 3 linearly independent components. Second, $\mathcal{T}$ is invariant under the flow of the Hamiltonian (7), as Eqs. (3) and (6) readily imply [^1] This equation can be obtained by a direct combination of Eqs. (6) and (7), or from the fact that Eqs. (1) and (4) imply $\nabla_{u}C^{a}=0$, which is a rewriting of (13) given that $\{C^{\alpha},H\}=\mathrm{d}C^{\alpha}/\mathrm{d}\bar{\tau}$.
$\{C^{\alpha},H\}=-\Gamma^{\alpha}_{\beta\gamma}p^{\beta}C^{\gamma}\,,$ (13)